OpenStudy (flowey15):

What is the radius of a circle with the equation x^2 + y^2 – 14x + 10y = 250? A) 9 units B) 12 units C) 15 units D) 18 units

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

do you know how to complete the square? you have to do it twice here, once for x and once for y

OpenStudy (flowey15):

no, I dont

OpenStudy (misty1212):

then there is no way for you to solve this you have to make it look like \[(x-h)^2+(y-k)^2=r^2\] to find \(r\)

OpenStudy (flowey15):

I know that part..i just don't know how

OpenStudy (misty1212):

i can show you if you like, it takes a few steps but if you have never seen completing the squarer, it might now make any sense

OpenStudy (misty1212):

start with \[x^2-14x+y^2+10y=250\] i.e group the terms together q

OpenStudy (misty1212):

now for the x what is half of 14?

OpenStudy (flowey15):

7

OpenStudy (misty1212):

and what is \(7^2\)?

OpenStudy (flowey15):

49, think

OpenStudy (misty1212):

yes so first step is to turn \[x^2-14x+y^2+10y=250\] in to \[(x-7)^2+y^2+10y=250+49\]

OpenStudy (misty1212):

now we do the same for y what is half of 10?

OpenStudy (dinamix):